# An effective control of large systems of active particles: An application to evacuation problem

### v2

@@ Abstract
[tv] organization=Research and Development Centre, TerraVox Global Ltd, city=Paphos, postcode=8200, country=Cyprus
[…]
@@ 1 Introduction
The manipulation of large systems of active particles, especially controlling their collective behavior, has become a fundamental problem in recent decades, initiated by emerging new areas of application. The examples, across diverse domains, include crowd management [1, 2], controlling of robotic swarms [3], coordinated material transport [4], etc. The concept of active matter [5] offers an appropriate framework for modeling such ensembles, comprised of agents that consume energy to move, interact with their environment, and adjust their direction of motion subject to external signals [6, 7, 8, 9]. Synthetic microswimmers, swarming robots, colonies of bacteria, fish schools, flocks of birds, groups of humans [10, 9] – all these systems of living and non-living agents represent an active matter.
@@
The primary goals of controlling active particle systems include navigating them in complex environments and supporting synchronized collective behavior [11, 12, 13]. This can be addressed by classical optimal control theory, which requires a complete knowledge of the environment and dynamics (similar to Zermelo’s navigation problem [14]), or by reinforcement learning (RL), where an agent learns strategies through trial-and-error [15, 16]. The latter is particularly applicable when an agent has only partial, local information or when the impact of noise is significant. A key limitation of such control problems is that their solution relies on the individual manipulation of each participating particle. For ensembles comprising hundreds of agents a straightforward application of RL, with individual-level control, becomes computationally intractable.
@@
To overcome this limitation, a common strategy involves guiding a group through a leader or multiple leaders, when active particles just follow the leader(s) [17, 18, 19, 20]. This leader-follower paradigm significantly simplifies the control problem by focusing strategies solely on the leader(s). Nevertheless, complexities persist, particularly when managing multiple groups of different size and location, or when groups cannot be manipulated simultaneously or possess non-coinciding goals [21, 22]. Among important examples of such challenging problems is an effective guidance of large groups of people, especially – their evacuation from hazardous places. Here a leader (rescuer) has the goal to evacuate all people guiding them to the exit(s) in shortest time. Given the inherent randomness and non-negligible noise within these systems, reinforcement learning (RL) seems to be the most suitable approach. However, a straightforward application of standard RL is neither computationally efficient nor effective in achieving the ultimate goal.
@@
To this end, we propose an application of auxiliary (artificial) “pseudo-gravitational” forces acting on the system, as a part of the environment; this helps the leader to find the most effective guidance strategy. Such a combination of RL with artificial forces [23] results in a very effective method to control large ensembles of active particles. To implement our new method for the evacuation problem, we utilize the generalized Vicsek model. This model takes into account not only interactions between active particles (as in the conventional Vicsek model [5]), but also between particles and the leader.
@@ 2.1 General statement of the evacuation problem
The problem of effective evacuation, especially in the case of panicking crowd is extremely important. Poor building planning or lack of awareness of escape routes in hazardous events (natural disasters, terrorist attacks, etc.) can lead to fatalities for many reasons, including stampedes [24, 25, 26, 27].
@@
Generally, evacuation studies employ several main methods [28]: indoor experiments with real people [29], numerical simulations (often incorporating machine learning for optimization) [24, 30, 31, 32], and hybrid approaches combining both [33, 34, 35]. The goal of such an analysis is to reveal some rules of crowd behavior, translate them into mathematical models, and reconstruct the crowd dynamics in numerical experiments. Machine learning can be used to optimize the realism or efficiency of evacuation.
Among the first studies of evacuation was Helbing’s work [1], where the social force model (SFM) has been pioneered. Based on visual patterns of people’s behavior, subject to a combination of physical interactions and the action of the social norm of keeping apart from unknown persons, the authors proposed an appropriate mathematical model. SFM was then successfully applied to describe panicking crowds [32]. This model is very general, and may be easily adopted, by respective modifications, to various special cases. SFM-like models are able to reproduce various self-organizing phenomena, such as lane formation, clogging at exits [36, 37, 38] customer dynamics [39], etc.
The main limitation of such models is the computational complexity, associated with the complete implementation of the SFM for a large system of agents. Therefore the lattice-automata approach, was widely used, with the set of discrete actions and states. Here the pedestrians are individual ”particles”, with an appropriate speed and size; their movement is governed by spatial transition probabilities. This allows encoding of simple rules of behavior in numerical simulations, which is realized in a number of available software ( Legion [40], EXITUS [30], Simulex [41], EXIT89 [42], PEDFLOW [43], etc.)
Further development of evacuation models has continued for the last two decades, and numerous new models have been suggested. These include models that allow evacuation testing of a building using the respective simulations at the planning stage [44]. Similarly, there are several studies, where various evacuation strategies can be tested for buildings already built [45, 46, 47, 48, 49].
@@
Various methods and models have been proposed to solve this problem. For example, a Markov Decision Process with intermediate decision-points has been employed [49, 39], incorporating the Social Force Model (SFM) [1] within a discrete state-action space. The problem has also been framed as multi-agent learning [34], treating groups of evacuating people as agents. To realistically describe inter-person interactions, SFM modifications have been explored, such as attributing group formation to attractive interactions. A two-level hierarchical approach, combining macro-level path programming with micro-level pedestrian-obstacle and inter-pedestrian interactions, was used in [48].
In the present study, we analyze the evacuation by an informed rescuer-leader, which guides the pedestrians towards the exit. We explore the conventional Viscek model [5] in the context of pedestrian motion dynamics and propose the ”generalized Viscek model” to account for the leader. Futher, we apply the method of reinforcement learning (RL) [15] to train the leader’s policy. The leader is supposed to be a robot which acts according to an optimal strategy, found in a learning process. The idea of a robot-assistant evacuation seems to be rather fruitful, see e.g. a review [50]. Further, we develop this approach and describe the problem setting.
@@ 2.2 Generalized Vicsek model
Here we introduce a generalized Vicsek model, which describes motion of people, guided by a leader. Generally, a motion of a person surrounded by a crowd may be decomposed into two parts – the one depending only on the person itself, and the other, imposed by the crowd. Often this results in self-organized patterns. One of the simplest models quantifying self-organization of moving agents, is the celebrated Vicsek model [5]. In spite of its simplicity, it demonstrates a rich behavior, with a large potential for further extensions, see e.g. [51, 52, 53, 54]. The model is based on a few simple rules and takes into account random fluctuations. In the original version [5], the main rule states, that each time step a moving agent adopts its direction of motion, to the average direction of all neighbours located within the distance $r$ from the agent. The absolute velocities of the agents are kept constant and the system is subjected to a white noise. Mathematically, the Vicsek model is formulated for a two-dimensional motion as following:
@@
The conventional Vicsek model reflects an important feature of a crowd – the so-called ”herd mentality” [55]. It reflects the empirical observation that the individuals may be influenced by their peers and adopt their behavior. The main difference between the current and standard Vicsek model is the introduction of a leader with an independent action policy. While in the original Vicsek model all individuals are peers, the motion of an individual in our model is affected by both – by the motion of the crowd and that of a particular individual, called leader. The present model assumes a single leader, although it may be straightforwardly generalized for multiple leaders. The leader is characterized by a stronger influence on the crowd, as compared to a common individual. Namely, an individual will preferably follow the direction of a leader and not of the neighbors. This extension is essential in the evacuation problem, where the leader is responsible for guiding individuals to the exit. To take into account an impact of a leader on a motion of a person, we generalize accordingly Vicsek model. We assume that the leader possesses its influence zone – a circle of radius $r_{\rm L}$, and its influence is characterized by the coefficient $q\in[0,1]$ – the enslaving degree, which quantifies the relative influence of the leader and neighbors. Hence we formulate the generalized Vicsek model as following [^1] Note that Eq. (2) is written in the original notations of Vicsek model, so that the expression $q\,\theta_{l}(t)+(1-q)\,\theta_{\rm v,i}(t)$ denotes the weighted average direction of the two directions with angles $\theta_{l}(t)$ and $\theta_{\rm v,i}(t)$ and corresponding weights $q$ and 1q1-q. This averaging is performed according to the procedure described in B [see Eqs. (23), and (24)]..
@@ 2.3 Evacuation simulation
is the unit vector for $i$-th particle to the exit destination. At the end, having reached the evacuation destination – the small circle with an exit point at the center, Figure 2, an individual is considered saved. For clear simulations and training, the environment has been implemented with the interface of OpenAI Gym environment [^2] Source is available on GitHub: https://github.com/cinemere/evacuation with the interface of OpenAI Gym environment (https://gymnasium.farama.org)[56]..
@@
The reward is used to train the leader in the evacuation environment, to achieve the best action policy. Here the main goal is to reduce the total evacuation time of the individuals, as much as possible. To achieve this goal an appropriate reward is to be tailored. The straightforward application of RL implies that a reward is given only after achieving the goal. This however results in a sparse rewarding, which is computationally very challenging, therefore we apply an approach of intermediate reward for low-level goals, similar to Ref. [57], and give positive reinforcement for each individual entering the exit zone. In Algorithm 1 we describe the procedure of reward computation. Here, $R_{t}$ denotes the reward at time step t; $N_{\texttt{exiting}}^{\texttt{new}}(t)$ is the number of individuals who just entered the Exit Zone at time step t; $\texttt{t}_{\texttt{max}}$ is the maximum episode length; $w_{A}=15$ and $w_{B}=10$ are empirical parameters influencing the amount of positive reinforcement; and $p=1$ is the step penalty for the time spent.
@@ 3 Methods
The idea to exploit potential-field methods to guide an agent was very fruitful in robotics, see e.g. [58, 59, 60, 61]. Here the goals and obstacles are treated as artificial ”fields” that exert virtual forces on a robot, helping in this way its navigation. In other words, a goal creates an ”attractive” basin, while obstacles create ”repulsive” hills, and the agent moves under the resulting force $\nabla U$. Because only local gradients are needed, the method yields fast, reactive navigation. Analogous ideas appear in reinforcement learning – by adding a “potential” term to the reward, one shapes a value landscape that nudges the learner in desirable directions while preserving the optimal policy. Hence artificial potential fields give an elegant, physically motivated way to steer agents. Inspired by these ideas we propose artificial force fields, which are also gradients of the respective artificial potentials. Since the potentials are expressed in terms of the inverse power of the according distances, we call them “pseudo-gravitational”. Note, however, that the nature of our potentials differ from this of standard artificial potentials, since they include the number of active particles in a group; below we present their detailed definition.
@@
There are different ways of defining the policy. For instance, the policy can be defined through a feed-forward neural network that straightforwardly processes the full set of individual positions as its input (note, that in our setting the state space is continuous and infinite). Such an approach may be, however, prohibitively expensive, especially with respect to RL optimization. [^3] This approach either assumes that the high-dimensional state can be directly compressed into a lower-dimensional latent representation, or requires scaling the network width proportionally to the input dimension. Both options make training inefficient or computationally demanding. Alternatively, since all individuals are equivalent, it would be more beneficial to deal with some appropriate collective characteristics. This suggests an implementation of the Deep Sets encoding [62]. Here we deploy this approach with a specific encoding possessing a physical interpretation, which may be called “pseudo-gravitational” encoding. Note, that such a reduction is particularly important for large ensembles. Indeed, a straightforward RL approach would require processing of the full state of all individuals; this becomes intractable for sufficiently large ensembles.
@@ 3.2 Training details
In the present work, we compare three methods: feed-forward, transformer, and a novel approach based on pseudo-gravitational encoding. The primary distinction among these methods lies in their neural network architectures. It is important to note that the pseudo-gravitational encoding is not a standalone neural architecture, but rather a state-representation scheme that acts as the first layer of the network. The encoded forces are then passed to the same feed-forward network architecture as in the baseline model, with the only difference being the input representation. To train the action policy, we utilized the Robust Policy Optimization (RPO) [63] implementation from the CleanRL [64] package. The complete set of RPO agent hyperparameters is summarized in Table 4, while additional details on the RL algorithm are presented in A.
@@
To better capture interrelations between particles, we incorporated a learnable attention mechanism into ANN [65]. Here, information concerning the exit, leader, and pedestrians is concatenated into a single token-vector. In general, a token in a transformer is the smallest unit of information that the model treats as an independent entity. Each token has its own vector representation and participates in the self-attention mechanism, which learns relations between tokens. The choice of tokens is problem-dependent: in NLP they are words, in the computer vision they are image patches. In our case, it is natural to define tokens as pedestrians, the leader, and the exit, since these are the entities whose interrelations drive the evacuation dynamics. Formally, we represent each token as a tuple $T=(\vec{r},i)$, where $r$ is the position vector of the entity (relative to the leader, unless otherwise specified), and $i$ is a one-hot indicator of its type or zone. That is, we define:
@@
Attention is estimated by adding BERT-style [66] transformer block(s), which transforms representation of each token: $T\rightarrow T^{\prime}$ allowing the model to focus on relevant input data and enhancing computational efficiency. The feed-forward policy then takes these transformed tokens as input:

### v3

@@ 1 Introduction
The manipulation of large systems of active particles, especially controlling their collective behavior, has become a fundamental problem in recent decades, initiated by emerging new areas of application. The examples, across diverse domains, include crowd management [helbing1995social, Warren2018CollectiveMI], controlling of robotic swarms [swarm_robots1], coordinated material transport [Uspal2018TheoryOL], etc. The concept of active matter [VicsekINITIAL] offers an appropriate framework for modeling such ensembles, comprised of agents that consume energy to move, interact with their environment, and adjust their direction of motion subject to external signals [doi:10.1177/10597123231202593, OLCAY20191, BAR2024101701, Egor2021]. Synthetic microswimmers, swarming robots, colonies of bacteria, fish schools, flocks of birds, groups of humans [RevModPhys.88.045006, Egor2021] – all these systems of living and non-living agents represent an active matter.
@@
The primary goals of controlling active particle systems include navigating them in complex environments and supporting synchronized collective behavior [microsw1, microsw2, microsw3]. This can be addressed by classical optimal control theory, which requires a complete knowledge of the environment and dynamics (similar to Zermelo’s navigation problem [Zermelo]), or by reinforcement learning (RL), where an agent learns strategies through trial-and-error [RL, RL2]. The latter is particularly applicable when an agent has only partial, local information or when the impact of noise is significant. A key limitation of such control problems is that their solution relies on the individual manipulation of each participating particle. For ensembles comprising hundreds of agents a straightforward application of RL, with individual-level control, becomes computationally intractable.
@@
To overcome this limitation, a common strategy involves guiding a group through a leader or multiple leaders, when active particles just follow the leader(s) [Seyboth2014CooperativeCO, Nowzari2017EventtriggeredCA, Hu2021ADC, Long2019ACR]. This leader-follower paradigm significantly simplifies the control problem by focusing strategies solely on the leader(s). Nevertheless, complexities persist, particularly when managing multiple groups of different size and location, or when groups cannot be manipulated simultaneously or possess non-coinciding goals [Bernardi_2021, Albi2015InvisibleCO]. Among important examples of such challenging problems is an effective guidance of large groups of people, especially – their evacuation from hazardous places. Here a leader (rescuer) has the goal to evacuate all people guiding them to the exit(s) in shortest time. Given the inherent randomness and non-negligible noise within these systems, reinforcement learning (RL) seems to be the most suitable approach. However, a straightforward application of standard RL is neither computationally efficient nor effective in achieving the ultimate goal.
@@
To this end, we propose an application of auxiliary (artificial) “pseudo-gravitational” forces acting on the system, as a part of the environment; this helps the leader to find the most effective guidance strategy. Such a combination of RL with artificial forces [KALA2024465] results in a very effective method to control large ensembles of active particles. To implement our new method for the evacuation problem, we utilize the generalized Vicsek model. This model takes into account not only interactions between active particles (as in the conventional Vicsek model [VicsekINITIAL]), but also between particles and the leader.
@@ 2.1 General statement of the evacuation problem
The problem of effective evacuation, especially in the case of panicking crowd is extremely important. Poor building planning or lack of awareness of escape routes in hazardous events (natural disasters, terrorist attacks, etc.) can lead to fatalities for many reasons, including stampedes [Borve, Pradeepkumar, Peacock, Liu].
@@
Generally, evacuation studies employ several main methods [fang2022evacuation]: indoor experiments with real people [NATAPOV2022105483], numerical simulations (often incorporating machine learning for optimization) [Borve, 7332972, 8989970, 9442521], and hybrid approaches combining both [7373229, wang2019improved, 7152824]. The goal of such an analysis is to reveal some rules of crowd behavior, translate them into mathematical models, and reconstruct the crowd dynamics in numerical experiments. Machine learning can be used to optimize the realism or efficiency of evacuation.
Among the first studies of evacuation was Helbing’s work [helbing1995social], where the social force model (SFM) has been pioneered. Based on visual patterns of people’s behavior, subject to a combination of physical interactions and the action of the social norm of keeping apart from unknown persons, the authors proposed an appropriate mathematical model. SFM was then successfully applied to describe panicking crowds [9442521]. This model is very general, and may be easily adopted, by respective modifications, to various special cases. SFM-like models are able to reproduce various self-organizing phenomena, such as lane formation, clogging at exits [24a, Helbing2011, Bodrova2024] customer dynamics [tsukanov2021], etc.
The main limitation of such models is the computational complexity, associated with the complete implementation of the SFM for a large system of agents. Therefore the lattice-automata approach, was widely used, with the set of discrete actions and states. Here the pedestrians are individual ”particles”, with an appropriate speed and size; their movement is governed by spatial transition probabilities. This allows encoding of simple rules of behavior in numerical simulations, which is realized in a number of available software ( Legion [Legion], EXITUS [7332972], Simulex [THOMPSON1995149], EXIT89 [Fahy1994Exit8E], PEDFLOW [PEDFLOW], etc.)
Further development of evacuation models has continued for the last two decades, and numerous new models have been suggested. These include models that allow evacuation testing of a building using the respective simulations at the planning stage [SHIN2019545]. Similarly, there are several studies, where various evacuation strategies can be tested for buildings already built [Kuligowski2010, Kisko1985EVACNETAC, ABIR2022103083, zhang2021crowd, pugh2021deep].
@@
Various methods and models have been proposed to solve this problem. For example, a Markov Decision Process with intermediate decision-points has been employed [pugh2021deep, tsukanov2021], incorporating the Social Force Model (SFM) [helbing1995social] within a discrete state-action space. The problem has also been framed as multi-agent learning [wang2019improved], treating groups of evacuating people as agents. To realistically describe inter-person interactions, SFM modifications have been explored, such as attributing group formation to attractive interactions. A two-level hierarchical approach, combining macro-level path programming with micro-level pedestrian-obstacle and inter-pedestrian interactions, was used in [zhang2021crowd].
In the present study, we analyze the evacuation by an informed rescuer-leader, which guides the pedestrians towards the exit. We explore the conventional Viscek model [VicsekINITIAL] in the context of pedestrian motion dynamics and propose the ”generalized Viscek model” to account for the leader. Futher, we apply the method of reinforcement learning (RL) [RL] to train the leader’s policy. The leader is supposed to be a robot which acts according to an optimal strategy, found in a learning process. The idea of a robot-assistant evacuation seems to be rather fruitful, see e.g. a review [Roboteva]. Further, we develop this approach and describe the problem setting.
@@ 2.2 Generalized Vicsek model
Here we introduce a generalized Vicsek model, which describes motion of people, guided by a leader. Generally, a motion of a person surrounded by a crowd may be decomposed into two parts – the one depending only on the person itself, and the other, imposed by the crowd. Often this results in self-organized patterns. One of the simplest models quantifying self-organization of moving agents, is the celebrated Vicsek model [VicsekINITIAL]. In spite of its simplicity, it demonstrates a rich behavior, with a large potential for further extensions, see e.g. [PhysRevLett.75.4326, Gr_goire_2003, labasken, collective_turns]. The model is based on a few simple rules and takes into account random fluctuations. In the original version [VicsekINITIAL], the main rule states, that each time step a moving agent adopts its direction of motion, to the average direction of all neighbours located within the distance $r$ from the agent. The absolute velocities of the agents are kept constant and the system is subjected to a white noise. Mathematically, the Vicsek model is formulated for a two-dimensional motion as following:
@@
The conventional Vicsek model reflects an important feature of a crowd – the so-called ”herd mentality” [rast1963effects]. It reflects the empirical observation that the individuals may be influenced by their peers and adopt their behavior. The main difference between the current and standard Vicsek model is the introduction of a leader with an independent action policy. While in the original Vicsek model all individuals are peers, the motion of an individual in our model is affected by both – by the motion of the crowd and that of a particular individual, called leader. The present model assumes a single leader, although it may be straightforwardly generalized for multiple leaders. The leader is characterized by a stronger influence on the crowd, as compared to a common individual. Namely, an individual will preferably follow the direction of a leader and not of the neighbors. This extension is essential in the evacuation problem, where the leader is responsible for guiding individuals to the exit. To take into account an impact of a leader on a motion of a person, we generalize accordingly Vicsek model. We assume that the leader possesses its influence zone – a circle of radius $r_{\rm L}$, and its influence is characterized by the coefficient $q\in[0,1]$ – the enslaving degree, which quantifies the relative influence of the leader and neighbors. Hence we formulate the generalized Vicsek model as following [^1] Note that Eq. (2) is written in the original notations of Vicsek model, so that the expression $q\,\theta_{l}(t)+(1-q)\,\theta_{\rm v,i}(t)$ denotes the weighted average direction of the two directions with angles $\theta_{l}(t)$ and $\theta_{\rm v,i}(t)$ and corresponding weights $q$ and 1q1-q. This averaging is performed according to the procedure described in B [see Eqs. (23), and (24)]..
@@ 2.3 Evacuation simulation
is the unit vector for $i$-th particle to the exit destination. At the end, having reached the evacuation destination – the small circle with an exit point at the center, Figure 2, an individual is considered saved. For clear simulations and training, the environment has been implemented with the interface of OpenAI Gym environment [^2] Source is available on GitHub: https://github.com/cinemere/evacuation with the interface of OpenAI Gym environment (https://gymnasium.farama.org)[brockman2016openai]..
@@
The reward is used to train the leader in the evacuation environment, to achieve the best action policy. Here the main goal is to reduce the total evacuation time of the individuals, as much as possible. To achieve this goal an appropriate reward is to be tailored. The straightforward application of RL implies that a reward is given only after achieving the goal. This however results in a sparse rewarding, which is computationally very challenging, therefore we apply an approach of intermediate reward for low-level goals, similar to Ref. [dorigo1998], and give positive reinforcement for each individual entering the exit zone. In Algorithm 1 we describe the procedure of reward computation. Here, $R_{t}$ denotes the reward at time step t; $N_{\texttt{exiting}}^{\texttt{new}}(t)$ is the number of individuals who just entered the Exit Zone at time step t; $\texttt{t}_{\texttt{max}}$ is the maximum episode length; $w_{A}=15$ and $w_{B}=10$ are empirical parameters influencing the amount of positive reinforcement; and $p=1$ is the step penalty for the time spent.
@@ 3 Methods
The idea to exploit potential-field methods to guide an agent was very fruitful in robotics, see e.g. [Lynch, Khatib, Koren, Devlin]. Here the goals and obstacles are treated as artificial ”fields” that exert virtual forces on a robot, helping in this way its navigation. In other words, a goal creates an ”attractive” basin, while obstacles create ”repulsive” hills, and the agent moves under the resulting force $\nabla U$. Because only local gradients are needed, the method yields fast, reactive navigation. Analogous ideas appear in reinforcement learning – by adding a “potential” term to the reward, one shapes a value landscape that nudges the learner in desirable directions while preserving the optimal policy. Hence artificial potential fields give an elegant, physically motivated way to steer agents. Inspired by these ideas we propose artificial force fields, which are also gradients of the respective artificial potentials. Since the potentials are expressed in terms of the inverse power of the according distances, we call them “pseudo-gravitational”. Note, however, that the nature of our potentials differ from this of standard artificial potentials, since they include the number of active particles in a group; below we present their detailed definition.
@@
There are different ways of defining the policy. For instance, the policy can be defined through a feed-forward neural network that straightforwardly processes the full set of individual positions as its input (note, that in our setting the state space is continuous and infinite). Such an approach may be, however, prohibitively expensive, especially with respect to RL optimization. [^3] This approach either assumes that the high-dimensional state can be directly compressed into a lower-dimensional latent representation, or requires scaling the network width proportionally to the input dimension. Both options make training inefficient or computationally demanding. Alternatively, since all individuals are equivalent, it would be more beneficial to deal with some appropriate collective characteristics. This suggests an implementation of the Deep Sets encoding [DeepSets]. Here we deploy this approach with a specific encoding possessing a physical interpretation, which may be called “pseudo-gravitational” encoding. Note, that such a reduction is particularly important for large ensembles. Indeed, a straightforward RL approach would require processing of the full state of all individuals; this becomes intractable for sufficiently large ensembles.
@@ 3.2 Training details
In the present work, we compare three methods: feed-forward, transformer, and a novel approach based on pseudo-gravitational encoding. The primary distinction among these methods lies in their neural network architectures. It is important to note that the pseudo-gravitational encoding is not a standalone neural architecture, but rather a state-representation scheme that acts as the first layer of the network. The encoded forces are then passed to the same feed-forward network architecture as in the baseline model, with the only difference being the input representation. To train the action policy, we utilized the Robust Policy Optimization (RPO) [rahman2022robust] implementation from the CleanRL [huang2022cleanrl] package. The complete set of RPO agent hyperparameters is summarized in Table 4, while additional details on the RL algorithm are presented in A.
@@
To better capture interrelations between particles, we incorporated a learnable attention mechanism into ANN [vaswani2017attention]. Here, information concerning the exit, leader, and pedestrians is concatenated into a single token-vector. In general, a token in a transformer is the smallest unit of information that the model treats as an independent entity. Each token has its own vector representation and participates in the self-attention mechanism, which learns relations between tokens. The choice of tokens is problem-dependent: in NLP they are words, in the computer vision they are image patches. In our case, it is natural to define tokens as pedestrians, the leader, and the exit, since these are the entities whose interrelations drive the evacuation dynamics. Formally, we represent each token as a tuple $T=(\vec{r},i)$, where $r$ is the position vector of the entity (relative to the leader, unless otherwise specified), and $i$ is a one-hot indicator of its type or zone. That is, we define:
@@
Attention is estimated by adding BERT-style [kenton2019bert] transformer block(s), which transforms representation of each token: $T\rightarrow T^{\prime}$ allowing the model to focus on relevant input data and enhancing computational efficiency. The feed-forward policy then takes these transformed tokens as input: